OpenStudy (anonymous):

help! "What is the period, in seconds, of a simple pendulum of length 2 meters? Use the gravitational constant g = 9.8 m/s^2 and round your answer to two decimal places.

OpenStudy (nincompoop):

what do you know?

OpenStudy (anonymous):

that the pendulum is 2 meters long

OpenStudy (nincompoop):

how do you solve for the period?

OpenStudy (nincompoop):

\[T = 2 \pi \sqrt{\frac{L}{g}}\] L length of pendulum g gravitational constant

OpenStudy (anonymous):

using a calculator i got 2.838 which i guess rounds to 2.84

OpenStudy (nincompoop):

for your answer to be a correct approximation you will need to make sure you have the proper significant figures 2.00 meters 9.80 m/s^2 that is the only way you will be able to justify to a hundredth degree precision

OpenStudy (nincompoop):

that seems correct

OpenStudy (nincompoop):

good day

OpenStudy (anonymous):

and good day to you, nincompoop!